OpenStudy (anonymous):

Let x = 1 and Δx = 0.01. Find Δy. (Round your answer to five decimal places.) f(x) = x/x^2 + 5 Please help

OpenStudy (aum):

\[ \Delta y = f(x + \Delta x) - f(x) = f(1 + 0.01) - f(1) = f(1.01) - f(1) \]